OpenStudy (anonymous):

If cos Θ = negative 4 over 7, what are the values of sin Θ and tan Θ?

OpenStudy (anonymous):

what do you mean by that are you trying to find and angle? or we are using -4/7 to find the angle of cos , sin, and tan

OpenStudy (anonymous):

It is Pythagorean Identity

OpenStudy (anonymous):

I can do most of it, but when the cos gets squared off, I don't understand how the fraction changes.

OpenStudy (anonymous):

\[\cos \theta = \frac{ adj }{ hyp} \rightarrow \cos \theta=\frac{ -4 }{ 7 }\] \[\sin \theta=\frac{ opp }{hyp } and \tan \theta=\frac{ opp }{ adj}\]

OpenStudy (anonymous):

Are you certain that there was not more to the problem such as limitations on the angle?

OpenStudy (anonymous):

Not that I know of.

OpenStudy (anonymous):

If the terminal side of the angle is in the second quadrant, then the sine will be positve and the tangent negative; if the terminal side is in the third quadrant than the signs of sine and tangent would be reversed.

OpenStudy (anonymous):

Cosine is negative in the second and third quadrants. There wasnt' something like\[90<\theta<180\]or\[180<\theta<270\]as part of the problem?

OpenStudy (anonymous):

No, not for this one, it's done like this sample one, Substitute 1 over 4 into the Pythagorean Identity sin2 θ + cos2 θ = 1 and solve for cos θ. sin2 θ + cos2 θ = 1 the fraction 1 over 4 squared+ cos2 θ = 1 1 over 15 + cos2 θ = 1 cos2 θ = 15 over 16 the square root of cosine squared theta equals plus or minus the square root of the fraction 15 over 16 cos θ = plus or minus the square root of 15 over 4 Once you know sin θ = 1 over 4 and cos θ = plus or minus the square root of 15 over 4 , you can find the value of tan θ. tan θ = sine of theta over cosine of theta tan θ = the fraction 1 over 4 over the fraction plus or minus the square root of 15 over 4 tan θ = 1 over 4 times 4 over plus or minus square root of 15 tan θ = 1 over plus or minus square root of 15 tan θ = plus or minus square root of 15 over 15

OpenStudy (anonymous):

OK, they seem to be accepting the plus or minus as part of the answer indicating the ambiguity I have been asking about. I'll make a drawing of the situation on the next post.

OpenStudy (anonymous):

There is also a third quadrant version of this also